Suppose that there are 20 people are interested in buying good x who had a reservation price of 100 shillings and the 15th person had a reservation price of 50 shillings what will the demand curve look like


1
Expert's answer
2021-06-30T13:48:10-0400

If there are 20 people are interested in buying good x who had a reservation price of 100 shillings and the 15th person had a reservation price of 50 shillings, then the demand function is:

"(Q - 20)\/(15 - 20) = (P - 100)\/(50 - 100),"

"Q - 20 = 0.1P - 10,"

Qd = 0.1P + 10.